# Spatial disaggregation of time series

### v1

@@ Appendix A Proof of Theorem 1
$\boldsymbol{\Sigma}_{Y}$ $=\mathbf{C}\mathbf{A}^{-1}(\boldsymbol{\Sigma}_{U}\otimes\mathbb{I}_{n})(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}$
Hence, the gradient of $\ln L(\boldsymbol{\theta})$ is:
@@
$\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\boldsymbol{\beta}}$ $=\left(\frac{\partial\boldsymbol{\mu}}{\partial\boldsymbol{\beta}}\right)^{\top}\boldsymbol{\Sigma}_{Y}^{-1}\mathbf{e}=(\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top})\boldsymbol{\Sigma}_{Y}^{-1}\mathbf{e}$ (24)
$\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\sigma^{2}}$ $=\frac{1}{2}\mbox{tr}\left(\boldsymbol{\Sigma}_{Y}^{-1}\frac{\partial\boldsymbol{\Sigma}_{Y}}{\partial\sigma^{2}}\right)-\frac{1}{2}\mathbf{e}^{\top}\frac{\partial\boldsymbol{\Sigma}_{Y}^{-1}}{\partial\sigma^{2}}\mathbf{e}$
$=\frac{T}{2\sigma^{2}}-\frac{1}{2\sigma^{2}}\mathbf{e}^{\top}\boldsymbol{\Sigma}_{Y}^{-1}\mathbf{e}$ (25)
$\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\phi_{1}}$ $=\frac{1}{2}\mbox{tr}\left(\boldsymbol{\Sigma}_{Y}^{-1}\frac{\partial\boldsymbol{\Sigma}_{Y}}{\partial\phi_{1}}\right)-\frac{1}{2}\mathbf{e}^{\top}\boldsymbol{\Sigma}_{Y}^{-1}\frac{\partial\boldsymbol{\Sigma}_{Y}}{\partial\phi_{1}}\boldsymbol{\Sigma}_{Y}^{-1}\mathbf{e}$ (26)
$\frac{\partial\boldsymbol{\Sigma}_{Y}}{\partial\phi_{1}}$ $=\mathbf{CA}^{-1}\left(\frac{\partial\mathbf{\Sigma}_{U}}{\partial\phi_{1}}\right)(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}$
$\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\rho}$ $=\frac{1}{2}\mbox{tr}\left(\boldsymbol{\Sigma}_{Y}^{-1}\frac{\partial\boldsymbol{\Sigma}_{Y}}{\partial\rho}\right)-\left(\frac{\partial\boldsymbol{\mu}}{\partial\rho}\right)^{\top}\boldsymbol{\Sigma}_{Y}^{-1}\mathbf{e}+\frac{1}{2}\mathbf{e}^{\top}\boldsymbol{\Sigma}_{Y}^{-1}\frac{\partial\boldsymbol{\Sigma}_{Y}}{\partial\rho}\boldsymbol{\Sigma}_{Y}^{-1}\mathbf{e}$ (27)
where:
$\mathbf{e}$ $=\mathbf{Y}_{a}-\boldsymbol{\mu}=\mathbf{CA}^{-1}\mathbf{U}=\mathbf{CA}^{-1}(\mathbf{R}\otimes\mathbb{I}_{n})\boldsymbol{\epsilon}=\mathbf{M}\boldsymbol{\epsilon}$
$\boldsymbol{\epsilon}$ $=(\epsilon_{11},\epsilon_{12},\cdots,\epsilon_{1T},\epsilon_{21},\cdots,\epsilon_{2T},\cdots,\epsilon_{nT})^{\top}$
$\mathbf{R}$ $=\begin{pmatrix}1&0&0&\cdots&0\\ \phi_{1}&1&0&\cdots&0\\ \phi_{1}^{2}&\phi_{1}&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \phi_{1}^{T-1}&\phi_{1}^{T-2}&\phi_{1}^{T-3}&\cdots&1\\ \end{pmatrix}$
Now, under Assumption 1, let $\mathbf{P}$ and $\mathbf{L}$ matrices of size $T\times T$, and $\mathbf{N}$ a matrix of size $1\times T$, $\gamma_{1}$ the skewness coefficient and $\gamma_{2}$ the kurtosis coefficient, it is clear that:
$\mathbb{E}(\mathbf{e})=\mathbf{0}$ (28)
$\mathbb{E}(\mathbf{e}\mathbf{e}^{\top})=\mathbb{E}(\mathbf{M}\boldsymbol{\epsilon}\boldsymbol{\epsilon}^{\top}\mathbf{M}^{\top})=\sigma^{2}\mathbf{M}\mathbf{M}^{\top}=\mathbf{C}\mathbf{\Sigma}_{Y}\mathbf{C}^{\top}$ (29)
$\mathbb{E}(\mathbf{e}^{\top}\mathbf{e})=\mbox{tr}(\mathbb{E}(\mathbf{ee}^{\top}))=\sigma^{2}\mbox{tr}(\mathbf{M}\mathbf{M}^{\top})=\sigma^{2}\sum_{t=1}^{T}(\mathbf{M}\mathbf{M}^{\top})_{tt}$ (30)
$\mathbb{E}(\mathbf{N}\mathbf{e}\mathbf{e}^{\top}\mathbf{P}\mathbf{e})=\sum_{i=1}^{n}\sum_{t=1}^{T}\mathbf{(NM)}_{it}(\mathbf{M}^{\top}\mathbf{PM})_{it,it}\mathbb{E}(\epsilon_{it}^{3})$
$=\sigma^{3}\gamma_{1}\sum_{i=1}^{n}\sum_{t=1}^{T}\mathbf{(NM)}_{it}(\mathbf{M}^{\top}\mathbf{PM})_{it,it}$ (31)
$\mathbb{E}(\mathbf{e}^{\top}\mathbf{L}\mathbf{e}\mathbf{e}^{\top}\mathbf{P}\mathbf{e})=\sum_{i=1}^{n}\sum_{t=1}^{T}(\mathbf{M}^{\top}\mathbf{LM})_{it,it}(\mathbf{M}^{\top}\mathbf{PM})_{it,it}\mathbb{E}(\epsilon_{it}^{4})+$
$\sum_{i\neq j}^{n}\sum_{t\neq k}^{T}(\mathbf{M}^{\top}\mathbf{LM})_{jk,it}(\mathbf{M}^{\top}\mathbf{PM})_{it,jk}\mathbb{E}(\epsilon_{it}^{2})\mathbb{E}(\epsilon_{jk}^{2})$
$\qquad=(3+\gamma_{2})\sigma^{4}\sum_{i=1}^{n}\sum_{t=1}^{T}(\mathbf{M}^{\top}\mathbf{LM})_{it,it}(\mathbf{M}^{\top}\mathbf{PM})_{it,it}+$
$\sigma^{4}\sum_{i\neq j}^{n}\sum_{t\neq k}^{T}(\mathbf{M}^{\top}\mathbf{LM})_{jk,it}(\mathbf{M}^{\top}\mathbf{PM})_{it,jk}$ (32)
Hence, using Equations (24)–(27) and (28)–(32) it is obtained that:
@@
$\mathbf{E}\left(\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\boldsymbol{\beta}}\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\boldsymbol{\beta}^{\top}}\right)$ $=(\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top})\boldsymbol{\Sigma}_{Y}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}$
@@
$=-\mathbf{E}\left(\frac{\partial^{2}\ln L(\boldsymbol{\theta})}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}\right)$ (33)
$\mathbf{E}\left(\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\boldsymbol{\beta}}\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\sigma^{2}}\right)=-\frac{1}{2}\gamma_{1}\sigma^{3}\sum_{i=1}^{n}\sum_{t=1}^{T}\mathbf{((\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top})\boldsymbol{\Sigma}_{Y}^{-1}M)}_{it}\times$
$(\mathbf{M}^{\top}\mathbf{\Sigma}_{Y}^{-1}\mathbf{M})_{it,it}=\mathcal{K}_{2}$ (34)
$\mathbf{E}\left(\frac{\partial^{2}\ln L(\boldsymbol{\theta})}{\partial\boldsymbol{\beta}\partial\sigma^{2}}\right)=\mathbf{E}\left((\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top})\boldsymbol{\Sigma}_{Y}^{-1}\mathbf{e}\right)=\mathbf{0}$
$\mathbf{E}\left(\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\boldsymbol{\beta}}\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\phi_{1}}\right)=-\frac{1}{2}\gamma_{1}\sigma^{3}\sum_{i=1}^{n}\sum_{t=1}^{T}\mathbf{((\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top})\boldsymbol{\Sigma}_{Y}^{-1}M)}_{it}\times$
@@
$(\mathbf{M}^{\top}\mathbf{\boldsymbol{\Sigma}_{Y}^{-1}\frac{\partial\boldsymbol{\Sigma}_{Y}}{\partial\phi_{1}}\boldsymbol{\Sigma}_{Y}^{-1}M})_{it,it}\mathcal{K}_{3}$ (35)

### v2

@@ Appendix B Proof of Theorem 1
    $\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}$ $=\mathbf{C}\mathbf{A}^{-1}(\boldsymbol{\Sigma}_{U}\otimes\mathbb{I}_{n})(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}$
    Hence, the gradient of $\ln L(\boldsymbol{\theta})$ with respect to $\boldsymbol{\beta}$ is:
@@
    $\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\boldsymbol{\beta}}$ $=\left(\frac{\partial\boldsymbol{\mu}_{a}}{\partial\boldsymbol{\beta}}\right)^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbb{E}=(\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top})\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{e}$ (33)
@@
    $\mathbb{E}\left(\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\boldsymbol{\beta}}\frac{\partial\ln L(\boldsymbol{\theta})}{\partial\boldsymbol{\beta}^{\top}}\right)$ $=(\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top})\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}$
@@
    $=-\mathbb{E}\left(\frac{\partial^{2}\ln L(\boldsymbol{\theta})}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}\right)$ (34)
    Under the model in Equation (2) and Assumptions 15, write the stacked observation equation as
    $\mathbf{Y}_{a}=\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\,\boldsymbol{\beta}+\mathbf{e},$
    with $\mathbb{E}(\mathbf{e})=\mathbf{0}$ and $\mathrm{Var}(\mathbf{e})=\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}$. The feasible GLS estimator for $\boldsymbol{\beta}$ is Bai et al. (2021):
    $\hat{\boldsymbol{\beta}}=\left[\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\widehat{\boldsymbol{\Sigma}}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\right]^{-1}\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\widehat{\boldsymbol{\Sigma}}_{\mathbf{Y}_{a}}^{-1}\mathbf{Y}_{a},$
    where $\widehat{\boldsymbol{\Sigma}}_{Y}$ is a estimator of $\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}$ obtained from the estimated nuisance parameters.
    When the true covariance $\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}$ is known, the estimator reduces to the infeasible GLS Bai et al. (2021):
    $\tilde{\boldsymbol{\beta}}$ $=\left[\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\right]^{-1}\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{Y}_{a}$
    $=\left[\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\right]^{-1}\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}(\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\,\boldsymbol{\beta}+\mathbf{e})$
    $=\left[\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\right]^{-1}\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\,\boldsymbol{\beta}+$
    $\left[\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\right]^{-1}\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{e}$
    $=\boldsymbol{\beta}+\left[\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\right]^{-1}\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{e}$
    Now decompose $\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta}$:
    $\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta}=\left[\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\right]^{-1}\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{e}.$
@@
    Under Assumption 1 and Assumptions 45, $\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-\frac{1}{2}}\mathbf{e}\sim\mathcal{N}_{T}(\mathbf{0},\mathbb{I}_{n})$, then
    $\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{e}\sim\mathcal{N}\big(\mathbf{0},\;\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\big).$
    Multiplying on the left by the inverse matrix yields
    $\tilde{\boldsymbol{\beta}}\sim\mathcal{N}_{k}\big(\boldsymbol{\beta},\;\big[\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\big]^{-1}\big).$
    Finally, to pass from the infeasible GLS $\tilde{\boldsymbol{\beta}}$ to the feasible estimator $\hat{\boldsymbol{\beta}}$ (FGLS), note that under Assumptions 13 and by Corollary 3.8 of Lehmann and Casella (1998, pg 448), $\widehat{\boldsymbol{\Sigma}}_{Y}$ is a consistent estimator of $\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}$, and the matrix
    $\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\widehat{\boldsymbol{\Sigma}}_{Y}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}$
    converges in probability to its probability limit
    $\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z},$
    which is nonsingular by Assumption 4. Let
    $\mathbf{F}_{T}=\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\widehat{\boldsymbol{\Sigma}}_{Y}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}$
    $\mathbf{X}_{T}=\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{e},$
    therefore by Theorem 2,
    $\hat{\boldsymbol{\beta}}\xrightarrow[T\to\infty]{d}\mathcal{N}_{k}\big(\boldsymbol{\beta},\;\big[\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\big]^{-1}\big).$
    Multiplying out the inverse in the displayed variance gives the variance matrix in Equation (18) of the theorem.
  3. 3.
    Using Theorem 8.22 of Lehmann and Casella (1998, pg 62), i) and ii) it is clear that iii) is satisfied.
Appendix C Proof of Theorem 2
Proof.
Let the naive predictor of $\mathbf{Y}$ be defined as
$\hat{\mathbf{Y}}^{1}=\frac{1}{n}\mathbf{C}^{\top}\mathbf{Y}_{a}=\mathbf{P}\mathbf{Y},\qquad\mathbf{P}=\mathbb{I}_{T}\otimes\frac{1}{n}\mathbf{J}_{n},$
where $\mathbf{J}_{n}=\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$ denotes the $n\times n$ matrix of ones and $\mathbb{I}_{T}$ the $T$-dimensional identity matrix. This estimator corresponds to the temporal vector of spatial means, obtained by averaging observations within each time period.
Since $\mathbf{Y}=\boldsymbol{\mu}+\mathbf{e}$ with $\mathbf{e}=\mathbf{A}^{-1}\mathbf{U}$, $\mathbb{E}(\mathbf{e})=\mathbf{0}$ and $\mathrm{Var}(\mathbf{e})=\mathbf{B}$, the expectation of $\hat{\mathbf{Y}}^{1}\mid\mathbf{Y}_{a}$ is
$\mathbb{E}(\hat{\mathbf{Y}}^{1}\mid\mathbf{Y}_{a})=\mathbf{P}(\boldsymbol{\mu}+\mathbf{BC}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}(\mathbf{Y}_{a}-\mathbf{C}\boldsymbol{\mu})),$
and using Equation (12), the bias of the predictor is therefore
$\mathrm{Bias}(\hat{\mathbf{Y}}^{1}\mid\mathbf{Y}_{a})$ $=\mathbb{E}(\hat{\mathbf{Y}}^{1}\mid\mathbf{Y}_{a})-\mathbb{E}(\mathbf{Y}\mid\mathbf{Y}_{a})$
$=\mathbf{P}(\boldsymbol{\mu}+\mathbf{BC}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}(\mathbf{Y}_{a}-\mathbf{C}\boldsymbol{\mu}))-\boldsymbol{\mu}-\mathbf{BC}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a})$
$=-(\mathbb{I}_{nT}-\mathbf{P})\left(\boldsymbol{\mu}+\mathbf{BC}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}(\mathbf{Y}_{a}-\mathbf{C}\boldsymbol{\mu})\right)$
$=-(\mathbb{I}_{nT}-\mathbf{P})\boldsymbol{\mu_{1a}}$ (35)
where $\boldsymbol{\mu}_{1a}=\boldsymbol{\mu}+\mathbf{BC}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}(\mathbf{Y}_{a}-\mathbf{C}\boldsymbol{\mu})$. The corresponding covariance matrix of the prediction error is
$\mathrm{Var}(\mathbf{Y}-\hat{\mathbf{Y}}^{1}\mid\mathbf{Y}_{a})$ $=(\mathbb{I}_{nT}-\mathbf{P})\,(\mathbf{B}-\mathbf{BC}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{CB})\,(\mathbb{I}_{nT}-\mathbf{P})$
$=(\mathbb{I}_{nT}-\mathbf{P})(\mathbb{I}_{nT}-\mathbf{P}_{1})\mathbf{B}(\mathbb{I}_{nT}-\mathbf{P})$ (36)
where $\mathbf{P}_{1}=\mathbf{BC}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}$. Furthermore, the matrix $\mathbf{A}^{-1}$ can be expressed as:
$\mathbf{A}^{-1}$ $=(\mathbb{I}_{T}\otimes(\mathbb{I}_{n}-\rho\mathbf{W}))^{-1}$
$=\mathbb{I}_{T}\otimes(\mathbb{I}_{n}-\rho\mathbf{W})^{-1},$
then, let Hence combining Equations (35) and (36), the mean squared error (MSE) of $\hat{\mathbf{Y}}^{1}$ with respect to $\mathbf{Y}$, and their expectations is given by:
$\mathrm{MSE}$ $(\hat{\mathbf{Y}}^{1},\mathbf{Y})=\frac{1}{nT}\sum_{i=1}^{nT}(\hat{Y}^{(1)}_{i}-Y_{i})^{2}$
$\mathbb{E}(\mathrm{MSE}$ $(\hat{\mathbf{Y}}^{1},\mathbf{Y})\mid\mathbf{Y}_{a})=\frac{1}{nT}\sum_{i=1}^{nT}\mathbb{E}((\hat{Y}^{(1)}_{i}-Y_{i})^{2}\mid\mathbf{Y}_{a})$
$=\frac{1}{nT}\sum_{i=1}^{nT}\mathbb{E}((\hat{Y}^{(1)}_{i}-Y_{i})^{2}\mid\mathbf{Y}_{a})$
$=\frac{1}{nT}\sum_{i=1}^{nT}\left(\mbox{Var}((\hat{Y}^{(1)}_{i}-Y_{i})\mid\mathbf{Y}_{a})+\mathbb{E}((\hat{Y}^{(1)}_{it}-Y_{it})\mid\mathbf{Y}_{a})^{2}\right)$
$=\frac{1}{nT}\left(\mbox{tr}(\mbox{Var}(\mathbf{Y}-\hat{\mathbf{Y}}^{1}\mid\mathbf{Y}_{a}))+\mbox{Bias}(\hat{\mathbf{Y}}^{1}\mid\mathbf{Y}_{a})\right)$
$=\|(\mathbb{I}_{nT}-\mathbf{P})\boldsymbol{\mu}_{1a}\|^{2}+\operatorname{tr}\big((\mathbb{I}_{nT}-\mathbf{P})(\mathbb{I}_{nT}-\mathbf{P}_{1})\mathbf{B}(\mathbb{I}_{nT}-\mathbf{P})\big)$
$=\|(\mathbb{I}_{nT}-\mathbf{P})\boldsymbol{\mu}_{1a}\|^{2}+\operatorname{tr}\big((\mathbb{I}_{nT}-\mathbf{P})(\mathbb{I}_{nT}-\mathbf{P}_{1})\mathbf{B}\big).$
The first term represents the squared bias due to ignoring the systematic component $\boldsymbol{\mu}$, while the second term accounts for the variance of the residual component.
For comparison, for the estimator proposed in Equation (15) $\mathbf{Y}$ given $\mathbf{Y}_{a}$, using property in Equation (12) and Theorem 1, it is clear that $\mathbb{E}(\hat{\mathbf{Y}}\mid\mathbf{Y}_{a})=\mathbf{Y}$, and their mean squared error is
$\mathbb{E}(\mathrm{MSE}(\hat{\mathbf{Y}},\mathbf{Y})\mid\mathbf{Y}_{a})$ $=\mbox{tr}(\mbox{Var}(\hat{\mathbf{Y}}\mid\mathbf{Y}_{a}))$
$=\operatorname{tr}\big(\mathbf{B}-\mathbf{B}\mathbf{C}^{\top}(\mathbf{C}\mathbf{B}\mathbf{C}^{\top})^{-1}\mathbf{C}\mathbf{B}+\mathbf{M}\mbox{Var}(\hat{\boldsymbol{\beta}})\mathbf{M}^{\top}\big)$
$=\operatorname{tr}((\mathbb{I}_{nT}-\mathbf{P}_{1})\mathbf{B})+\mbox{tr}(\mathbf{M}\mbox{Var}(\hat{\boldsymbol{\beta}})\mathbf{M}^{\top})$
where
$\mathbf{M}$ $=\mathbf{A}^{-1}\mathbf{Z}-\mathbf{B}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{CA}^{-1}\mathbf{Z}=(\mathbb{I}_{nT}-\mathbf{P}_{1})\mathbf{A}^{-1}\mathbf{Z}$
$=\left(\mathbb{I}_{T}\otimes[\mathbb{I}_{n}-m_{\rho}^{-1}\mathbf{F}^{-1}(\mathbf{F}^{-1})^{\top}\mathbf{J}_{n}]\right)\mathbf{A}^{-1}\mathbf{Z}=\mathbf{P}_{0}\tilde{\mathbf{Z}}$
$\mathbf{M}^{\top}\mathbf{M}$ $=\tilde{\mathbf{Z}}^{\top}\mathbf{P}_{0}^{\top}\mathbf{P}_{0}\tilde{\mathbf{Z}}$
From Theorem 1 we obtain that
[…]